OpenStudy (anonymous):

H-E-L-P! The pool company has developed new chemicals that transform the pH scale. Using the pH function p(t) = –log10 t as the parent function, explain which transformation would result in a y-intercept. Use complete sentences and show all translations on your graph. • p(t) + 1 • p(t + 1) • –1 • p(t)

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (campbell_st):

use this online graphing calculator to see what happens parent: y = -log(t) graph 1: y = -log(t) + 1 graph 2: y = -log(t + 1) graph 3: y = -1 x -log(t) https://www.desmos.com/calculator hope it helps....

OpenStudy (anonymous):

Thanks! I'll try it out, but how will I know which one will result in a 'y-intercept'?